OpenStudy (anonymous):

What is true about the solutions of a quadratic equation when the radicand in the quadratic formula is negative?

OpenStudy (anonymous):

No real solutions Two identical rational solutions Two different rational solutions Two irrational solutions

OpenStudy (anonymous):

i believe it is B

OpenStudy (anonymous):

LOL okay thx so much =)

OpenStudy (anonymous):

May i ask y you think that it's B?

OpenStudy (anonymous):

@mattsouth0015

OpenStudy (anonymous):

If we look at the quadratic formula: $$ ax^2 + bx + c = 0 \\ x_{1,2} = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a} $$Then the radicand will be the expression inside the radical which is \( (b^2 - 4ac) \) But if the radicant is negative then its square root \( \sqrt{b^2 - 4ac} \) cannot be a real number, because there is no real number that squares to be a negative number, so it will result in a complex number. That will make the solutions for \(x\) be complex as well and not real solutions. So the answer should be A

OpenStudy (anonymous):

Ugh! Oh well thx for the help @pitamar =)

OpenStudy (anonymous):

You're welcome =)